• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name mitchell

    Log entry time 16:11:38 on March29,2000

    Entry number 37203

    keyword=Day shift summary

    8:00 - 8:30 recycle Q3 after reset
    8:00 - 9:30 still problems at injector
    9:30 beam back but with funny tail in x-dist at C12 (from OTR)
    9:55 after adjustment x at OTR is OK
    9:55 take harp scans at target, : horizontal width is too narrow
    10:00 harp run 755 vert 145 by hor 100 : OK
    10:10 notice lots of motion at 1H01 (in prep for ep energy measurement) ask
    MCC to check
    10:20 high current checkout
    10:50 checkout complete, move carbon target in
    10:55 start ep energy measurement and 12C(e,e;p) calibration runs
    11:41 switch energy lock from C to A
    11:45 insert 4 dummy target for main experiment
    11:50 beam back but unstable - lots of motion, position feedback not working, injector problems
    2:15 beam is back !!
    2:20 start ep energy measurements and 10 dummy runs
    2:40 beam off - Mcc inserted viewer near lambertson and corrupted ep run !
    3:50 ep done , move target to he3 . we amke a recheck of diagnostics, OTR, harps
    4:00 MCC calls and reports that they cannot turn on fast raster

    All in all a prettyy bad shift